# Leveraging Enhanced Queries of Point Sets for Vectorized Map Construction

### v1

@@ Abstract
In autonomous driving, the high-definition (HD) map plays a crucial role in localization and planning. Recently, several methods have facilitated end-to-end online map construction in DETR-like frameworks. However, little attention has been paid to the potential capabilities of exploring the query mechanism. This paper introduces MapQR, an end-to-end method with an emphasis on enhancing query capabilities for constructing online vectorized maps. Although the map construction is essentially a point set prediction task, MapQR utilizes instance queries rather than point queries. These instance queries are scattered for the prediction of point sets and subsequently gathered for the final matching. This query design, called the scatter-and-gather query, shares content information in the same map element and avoids possible inconsistency of content information in point queries. We further exploit prior information to enhance an instance query by adding positional information embedded from their reference points. Together with a simple and effective improvement of a BEV encoder, the proposed MapQR achieves the best mean average precision (mAP) and maintains good efficiency on both nuScenes and Argoverse 2. In addition, integrating our query design into other models can boost their performance significantly. The code will be available at https://github.com/HXMap/MapQR.
1 Introduction
@@
Many existing works [13, 36] define map construction as a semantic segmentation task in the bird’s eye view (BEV) space, producing rasterized maps. Although they have achieved much success, they face limitations due to extensive post-processing required to acquire vectorized information. To overcome the limitations of segmentation-based methods, new approaches have arisen that predict point sets to construct maps, utilizing a DETR-like structure for end-to-end map construction [18, 15, 16, 23, 6, 31].
@@
DETR-like object detection methods predict a 4-dimensional bounding box information for each object by a set of learnable queries. As proved in DAB-DETR [17], each query in a decoder is composed of a decoder embedding (content information) and a learnable query (positional information). By contrast, HD map construction usually predicts a point set for each map element. In many state-of-the-art (SOTA) methods [15, 16, 6], point queries are used to probe information from BEV features, and each one predicts a point position. These predicted points are then grouped to form the detected map elements. Although information has been exchanged and enhanced in a self-attention module, it is still hard for a point query to contain all content information of the whole map element. In addition, point queries within the same map element may even have different semantic content information. We call this phenomenon as content conflicts. Therefore, previous point-query-based map construction methods have difficulty in learning desirable content information. In addition, current SOTA methods ignore the location information and simply use learnable queries that are randomly initialized.
To solve the aforementioned problem, we propose to use instance queries rather than point queries and add positional embedding, as shown in Fig. 1. Instead of predicting one position from each point query separately, we predict point positions simultaneously from each instance query to ensure consistency of content information in the same map element. To probe information from specific positions of BEV features, we generate positional embeddings from reference points like in Conditional DETR [20] (i.e., positional embeddings for each instance query). Then different positional embeddings will be added to each instance query, which becomes scattered queries. Therefore, each map element contains a group of scattered queries that share the same content part from a single instance query and different positional parts embedded from reference positions. The set of scattered queries is gathered back as an instance query to match a map element. We call this query a scatter-and-gather query. Since only instance queries are used as the input of a transformer decoder, it avoids content conflicts in the same map element. It also allows the proposed decoder to increase the number of queries to achieve higher accuracy without noticeably increasing the computational burden and memory usage. The query design is the fundamental cornerstone of our proposed method and, with a simple and effective improvement of a BEV encoder, constitutes our proposed MapQR.
@@ 2.1 Online Vectorized HD Map Construction
VectorMapNet [18] proposes the first end-to-end vectorized map learning framework and utilizes auto-regression to predict points sequentially. Later end-to-end methods for vectorized HD map construction become popular [15, 16, 25, 23, 6, 30, 31, 32, 33]. MapTR [15] treats online map construction as a point set prediction problem and designs a DETR-like framework [2], achieving state-of-the-art performance. Its improved version MapTRv2 [16] designs a decoupled self-attention decoder and auxiliary losses to improve the performance further. Instead of representing a map element by a point set, BeMapNet [23] uses piecewise Bézier curve, and PivotNet [6] uses pivot-based points for more elaborate modeling. Most of these works use point-based queries and should meet the problem stated in Sec. 1. Recently, StreamMapNet [32] utilizes multi-point attention for a wide perception range and exploits long-sequence temporal fusion. Unlike previous work, we provide an in-depth exploitation of the instance-based query and design scatter-and-gather query to probe both content and positional information accurately.
@@ 2.2 Multi-View Camera-to-BEV Transformation
HD map is commonly constructed under BEV view, and thus online HD map construction methods depend on converting visual features from camera perspective to BEV space. LSS [22] utilizes latent depth distribution to lift 2D image features to 3D space, and uses pooling later to aggregate to BEV features. BEVformer [14] depends on the transformer architecture [26] and designs spatial cross-attention based on visual projection. Unlike deformable attention [5] used in BEVformer, GKT [4] designs a geometry-guided kernel transformer. Besides, temporal information has also been explored to learn better BEV features [14, 21, 7]. All these methods can produce BEV features effectively, and our proposed decoder is compatible with all of them.
@@ 2.3 Detection Transformers
DETR [2] builds the first end-to-end object detector with transformer [26], eliminating the need for many hand-designed components. In the object detection task, it represents the object boxes as a set of queries, and directly uses the transformer encoder-decoder architecture to interact the queries with the image to predict the set of bounding boxes. Subsequent works [37, 37, 20, 28, 17, 12, 34] further optimize this paradigm and performance. In Deformable DETR [37], multi-scale features are incorporated to solve the problem of poor detection results of the vanilla DETR for small objects. In Conditional DETR [20] and DAB-DETR [17], the convergence of DETR is accelerated by adding positional embedding to learnable queries. In DN-DETR [12] and DINO [34], the actual boxes with noise are input into the decoder, so as to establish a more stable pairing relationship to accelerate convergence. The paradigm of DETR has been widely used in fields other than object detection tasks. For example, online vectorized HD map construction usually adopts DETR-like models [15]. Our model is also a DETR-like model. Leveraging point set representation of map elements, we design queries that can scatter and gather in a transformer decoder, and integrate positional information into the initialization of queries.
@@ 3.1 Overall Architecture
Giving surrounding images from a multi-camera rig as inputs, we first extract image features through a shared 2D backbone. These image features of the cameras’ perspective view (PV) are converted to BEV representations by feeding into a view transformation module, commonly referred to as BEV encoder. The resulted BEV features are denoted as , where , , represent the number of feature channels, height, and width, respectively. All mainstream 2D backbones [10, 19] and view transformation modules which can be used for PV-to-BEV transformation [4, 14] can be integrated into our model.
@@ 3.2 Decoder with Scatter-and-Gather Query
The detailed architectures of the proposed decoder with comparison to MapTR series [15, 16] are demonstrated in Fig. 3. MapTR [15] uses point queries as the input of decoder layers and calculates self-attention scores between them. This makes the computational complexity of self-attention much higher than our model because of much more involved queries.
By contrast, the proposed query design can ensure consistency of content in the same map element. Meanwhile, this design reduces the number of queries to reduce computational consumption in self-attention. Thus, it allows the involvement of more instance queries without significantly increasing memory consumption. Detailed experimental results are reported in Tab. 4.
@@
We have also tried adding positional embedding to instance query before self-attention modules, such as using learnable queries or encoding from the smallest bounding box covering the set of reference points. However, it did not achieve positive results, which are included in supplementary material. So, the positional embedding of the instance query is excluded in our model.
@@ 3.4 Matching Cost and Training Loss
where is the set of permutations with elements. is a pair-wise matching cost between predictions and GT map elements . Then we use matching of scattered instances to find an optimal point-to-point assignment between the set of scattered instances and the set of GT points. The matching is established by calculating the Manhattan distance, which is the same as in MapTR.
@@
[…]
Backbone |
Epoch |
[…]
| BeMapNet [23] |
R50 |
[…]
R50 |
SwinT |
[…]
| PivotNet [6] |
R50 |
[…]
SwinT |
[…]
| MapTR [15] |
R50 |
[…]
R50 |
SwinT |
SwinB |
[…]
| MapTR +SGQ |
R50 |
[…]
R50 |
SwinT |
SwinB |
[…]
| MapTRv2 [16] |
R50 |
[…]
R50 |
[…]
| MapTRv2 +SGQ |
R50 |
[…]
R50 |
[…]
| MapQR (Ours) |
R50 |
[…]
R50 |
SwinT |
SwinB |
@@ 4 Experiments
We compare the proposed method MapQR with SOTA methods, including MapTR series [15, 16], BeMapNet [23] and PivotNet [6]. To validate the effectiveness of the proposed SGQ decoder separately, we also integrate it into the MapTR series by replacing their point-query-based decoders with the proposed decoder.
@@ 4.1 Settings
Datasets. We evaluate our method on popular self-driving datasets nuScenes [1] and Argoverse 2 [29]. These datasets provide surrounding images captured by self-driving cars in hundreds of city scenes. Each sample contains RGB images in nuScenes and RGB images in Argoverse 2. nuScenes provides 2D vectorized map elements, and Argoverse 2 provides 3D vectorized map elements as ground truth.
Evaluation Metric. Following previous works [13, 15], three static map categories are used for performance evaluation, i.e., lane divider, pedestrian crossing, and road boundary. Chamfer distance is used to determine whether the prediction matches the ground truth. Besides the thresholds used in MapTR [15], the results are also evaluated with smaller thresholds used in BeMapNet [23]. The corresponding mAP metrics are denoted as mAP and mAP, respectively.
Training and Inference Details. We follow most settings as in the MapTR series [15, 16], and the modifications will be emphasized. The BEV feature is set to be to perceive [] for rear to front, and [] for left to right. We use instance queries to detect map element instances, and each instance is modeled by sequential points. (We set and .)
@@ 4.2 Comparisons with State-of-the-art Methods
Results on nuScenes. Following the previous experiment settings, 2D vectorized map elements are predicted by different methods. The experimental results are listed in Tab. 1. It can be seen that the proposed MapQR outperforms all other SOTA methods by a large margin for both mAP and mAP, under the same settings (i.e., backbone and training epoch). Compared with MapTR and MapTRv2, BeMapNet [23] and PivotNet [6] get better results under smaller thresholds (i.e., mAP) thanks to their elaborate modeling. Still, the proposed method without elaborate modeling for map elements can outperform all of them. It implies that our instance query and improved encoder benefit for more accurate prediction. The inference speed of our method is about FPS, which meets the efficiency requirements of many scenarios.
To validate the effectiveness of the proposed SGQ decoder, it is integrated into both MapTR [15] and MapTRv2 [16], and all other settings except for the decoder are kept unchanged. The corresponding methods are denoted as MapTR+SGQ and MapTRv2+SGQ, respectively. Their performance and relative improvements are also listed in Tab. 1. For MapTR+SGQ, when trained with ResNet50 for epochs, our decoder helps to improve over mAP for both thresholds. Under other settings, the SGQ decoder can also enhance MapTR significantly. While MapTRv2 [16] uses decoupled self-attention in the decoder and some auxiliary supervision to improve results, simply replacing their decoder with ours can still step further, especially for the evaluation under stricter thresholds. For inference speed, introducing our decoder slows down only a little compared with original MapTR or MapTRv2.
@@
[…]
| MapTR [15] | |||
| MapTRv2 [16] | |||
[…]
Results on Argoverse 2. In Tab. 2, we provide the experimental results on Argoverse 2. These testings are all trained epochs with ResNet50 as the backbone. Since Argoverse 2 provides 3D vectorized map elements as ground truth, 3D map elements can be predicted directly () as in MapTRv2. For both 2D and 3D prediction, the proposed MapQR achieves the best performance, especially for stricter thresholds (mAP). More detailed experimental results are included in the supplementary material.

### v2

# 11(2.5, -0.1)

@@ Abstract
In autonomous driving, the high-definition (HD) map plays a crucial role in localization and planning. Recently, several methods have facilitated end-to-end online map construction in DETR-like frameworks. However, little attention has been paid to the potential capabilities of exploring the query mechanism for map elements. This paper introduces MapQR, an end-to-end method with an emphasis on enhancing query capabilities for constructing online vectorized maps. To probe desirable information efficiently, MapQR utilizes a novel query design, called scatter-and-gather query, which is modelled by separate content and position parts explicitly. The base map instance queries are scattered to different reference points and added with positional embeddings to probe information from BEV features. Then these scatted queries are gathered back to enhance information within each map instance. Together with a simple and effective improvement of a BEV encoder, the proposed MapQR achieves the best mean average precision (mAP) and maintains good efficiency on both nuScenes and Argoverse 2. In addition, integrating our query design into other models can boost their performance significantly. The source code is available at https://github.com/HXMap/MapQR.
Keywords:
Autonomous driving Bird’s-eye-view (BEV) Vectorized map construction DETR1 Introduction
@@
Many existing works [12, 35] define map construction as a semantic segmentation task in the bird’s eye view (BEV) space, producing rasterized maps. They face limitations due to extensive post-processing required to acquire vectorized information. To overcome the limitations of segmentation-based methods, new approaches have arisen that predict point sets to construct maps, utilizing a DETR-like structure for end-to-end map construction [18, 14, 15, 22, 6, 30, 17, 36].
@@
DETR-like object detection methods predict a 4-dimensional bounding box information for each object by a set of learnable queries. As proved in DAB-DETR [16], each query in a decoder is composed of a decoder embedding (content information) and a learnable query (positional information). By contrast, HD map construction usually predicts a point set for each map instance. In many state-of-the-art (SOTA) methods [14, 15, 6], point queries are used to probe information from BEV features, and each one predicts a point position. These predicted points are finally grouped to form the detected map elements. As a result, the relation among point queries for the same element lacks explicit modelling, which would impede efficient learning for desirable content information. In addition, current SOTA methods ignore the explicit location information and simply use learnable queries that are randomly initialized.
To solve the aforementioned problems, we propose to use instance queries rather than point queries and add positional embedding, as shown in Fig. 1. Instead of predicting one position from each point query separately, we predict point positions simultaneously from each instance query. To probe information from specific positions of BEV features, we generate positional embeddings from reference points like in Conditional DETR [20] (i.e., positional embeddings for each instance query). Then different positional embeddings will be added to one instance query, which becomes scattered queries for each instance. Therefore, each map element contains a group of scattered queries that share the same content part from a single instance query and different positional parts embedded from reference positions. The set of scattered queries is gathered back as an instance query to enhance information within each instance. We call this query a scatter-and-gather query. The relations within a single instance is explicitly modelled as shared content and different positions, which are enhanced by scattering and gathering. It also allows the proposed decoder to increase the number of queries to achieve higher accuracy without noticeably increasing the computational burden and memory usage. The query design is the fundamental cornerstone of our proposed method and, with a simple and effective improvement of a BEV encoder, constitutes our proposed MapQR.
@@ 2.1 Online Vectorized HD Map Construction
VectorMapNet [18] proposes the first end-to-end vectorized map learning framework and utilizes auto-regression to predict points sequentially. Later end-to-end methods for vectorized HD map construction become popular [14, 15, 24, 22, 6, 29, 30, 31, 32, 17, 36]. MapTR [14] treats online map construction as a point set prediction problem and designs a DETR-like framework [2], achieving state-of-the-art performance. Its improved version MapTRv2 [15] designs a decoupled self-attention decoder and auxiliary losses to improve the performance further. Instead of representing a map element by a point set, BeMapNet [22] uses piecewise Bézier curve, and PivotNet [6] uses pivot-based points for more elaborate modeling. Most of these works use point-based queries and should meet the problem stated in Sec. 1. Recently, StreamMapNet [31] utilizes multi-point attention for a wide perception range and exploits long-sequence temporal fusion. Unlike previous work, we provide an in-depth exploitation of the instance-based query and design scatter-and-gather query to probe both content and positional information accurately.
@@ 2.2 Multi-View Camera-to-BEV Transformation
HD map is commonly constructed under BEV view, and thus online HD map construction methods depend on converting visual features from camera perspective to BEV space. LSS [21] utilizes latent depth distribution to lift 2D image features to 3D space, and uses pooling later to aggregate to BEV features. BEVformer [13] depends on the transformer architecture [25] and designs spatial cross-attention based on visual projection. Unlike deformable attention [5] used in BEVformer, GKT [4] designs a geometry-guided kernel transformer. All these methods can produce BEV features effectively, and our proposed decoder is compatible with all of them.
@@ 2.3 Detection Transformers
DETR [2] builds the first end-to-end object detector with transformer [25], eliminating the need for many hand-designed components. In the object detection task, it represents the object boxes as a set of queries, and directly uses the transformer encoder-decoder architecture to interact the queries with the image to predict the set of bounding boxes. Subsequent works [37, 37, 20, 27, 16, 11, 33] further optimize this paradigm and performance. In Deformable DETR [37], multi-scale features are incorporated to solve the problem of poor detection results of the vanilla DETR for small objects. In Conditional DETR [20] and DAB-DETR [16], the convergence of DETR is accelerated by adding positional embedding to learnable queries. In DN-DETR [11] and DINO [33], the actual boxes with noise are input into the decoder, so as to establish a more stable pairing relationship to accelerate convergence. The model in our method is a DETR-like model. Leveraging point set representation of map elements, we design queries that can scatter and gather in a transformer decoder, and integrate positional information into the initialization of queries.
@@ 3.1 Overall Architecture
Given surrounding images from a multi-camera rig as inputs, we first extract image features through a shared 2D backbone. These image features of the cameras’ perspective view (PV) are converted to BEV representations by feeding into a view transformation module, commonly referred to as BEV encoder. The resulted BEV features are denoted as , where , , represent the number of feature channels, height, and width, respectively. All mainstream 2D backbones [9, 19] and view transformation modules which can be used for PV-to-BEV transformation [4, 13] can be integrated into our model.
@@ 3.2 Decoder with Scatter-and-Gather Query
The detailed architectures of the proposed decoder with comparison to MapTR series [14, 15] are demonstrated in Fig. 3. MapTR [14] uses point queries as the input of decoder layers and calculates self-attention scores between them. This makes the computational complexity of self-attention much higher than our model because of much more involved queries. By contrast, the proposed query design can enhance information in the same map element. Meanwhile, this design reduces the number of queries to reduce computational consumption in self-attention. Thus, it allows the involvement of more instance queries without significantly increasing memory consumption. Detailed experimental results are reported in Tab. 6.
@@
We have also tried adding positional embedding to instance query before self-attention modules, such as using learnable queries or encoding from the smallest bounding box covering the set of reference points. However, it did not achieve positive results, which are included in the supplementary material. So, the positional embedding of the instance query is excluded in our model.
@@ 3.4 Matching Cost and Training Loss
where is the set of permutations with elements. is a pair-wise matching cost between predictions and GT map elements . Then we use matching of scattered instances to find an optimal point-to-point assignment between the set of scattered instances and the set of GT points. The matching is established by calculating the Manhattan distance.
@@ 4 Experiments
We compare the proposed method MapQR with SOTA methods, including MapTR series [14, 15], BeMapNet [22], PivotNet [6] and StreamMapNet [31]. To validate the effectiveness of the proposed SGQ decoder separately, we also integrate it into the MapTR series and StreamMapNet by replacing their point-query-based decoders with the proposed decoder.
@@ 4.1 Settings
Datasets. We evaluate our method on public datasets nuScenes [1] and Argoverse 2 [28]. These datasets provide surrounding images captured by self-driving cars in hundreds of city scenes. Each sample contains RGB images in nuScenes and RGB images in Argoverse 2. nuScenes provides 2D vectorized map elements, and Argoverse 2 provides 3D vectorized map elements as ground truth.
Evaluation Metric. Following previous works [12, 14], three static map categories are used for performance evaluation, i.e., lane divider, pedestrian crossing, and road boundary. Chamfer distance is used to determine whether the prediction matches the ground truth. Besides the thresholds used in MapTR [14], the results are also evaluated with smaller thresholds used in BeMapNet [22]. The corresponding mAP metrics are denoted as mAP2 and mAP1, respectively.
Method | Backbone | Epoch | APdiv | APped | APbou | mAP1 | FPS |
BeMapNet [22] | R50 | 30 | 46.9 | 39.0 | 37.8 | 41.3 | 8.3 |
R50 | 110 | 52.7 | 44.5 | 44.2 | 47.1 | 8.3 | |
SwinT | 30 | 49.1 | 42.2 | 39.9 | 43.7 | 7.9 | |
PivotNet [6] | R50 | 24 | 41.4 | 34.3 | 39.8 | 38.5 | - |
SwinT | 24 | 45.0 | 36.2 | 41.2 | 40.8 | - | |
MapTR [14] | R50 | 24 | 30.7 | 23.2 | 28.2 | 27.3 | 24.2 |
R50 | 110 | 40.5 | 31.4 | 35.5 | 35.8 | 24.2 | |
swinT | 24 | 30.4 | 24.5 | 29.9 | 28.9 | 18.2 | |
SwinB | 24 | 36.5 | 29.0 | 35.2 | 33.5 | 9.3 | |
MapTRv2 [15] | R50 | 24 | 40.0 | 35.4 | 36.3 | 37.2 | 19.6 |
R50 | 110 | 49.0 | 43.6 | 43.7 | 45.4 | 19.6 | |
StreamMapNet [31] | R50 | 24 | 42.9 | 32.3 | 33.2 | 36.2 | 19.9 |
MapQR(Ours) | R50 | 24 | 49.9 | 38.6 | 41.5 | 43.3 | 17.9 |
R50 | 110 | 57.3 | 46.2 | 48.1 | 50.5 | 17.9 | |
swinT | 24 | 49.3 | 38.5 | 41.1 | 43.0 | 14.8 | |
swinB | 24 | 52.3 | 41.4 | 43.2 | 45.6 | 8.3 |
Method | Backbone | Epoch | APdiv | APped | APbou | mAP2 | FPS |
BeMapNet[22] | R50 | 30 | 62.3 | 57.7 | 59.4 | 59.8 | 8.3 |
R50 | 110 | 66.7 | 62.6 | 65.1 | 64.8 | 8.3 | |
SwinT | 30 | 64.4 | 61.2 | 61.7 | 62.4 | 7.9 | |
PivotNet [6] | R50 | 24 | 56.5 | 56.2 | 60.1 | 57.6 | - |
SwinT | 24 | 60.6 | 59.2 | 62.2 | 60.6 | - | |
MapTR [14] | R50 | 24 | 51.5 | 46.3 | 53.1 | 50.3 | 24.2 |
R50 | 110 | 59.8 | 56.2 | 60.1 | 58.7 | 24.2 | |
swinT | 24 | 45.2 | 52.7 | 52.3 | 50.1 | 18.2 | |
SwinB | 24 | 58.7 | 50.6 | 58.4 | 55.9 | 9.3 | |
MapTRv2 [15] | R50 | 24 | 62.4 | 59.8 | 62.4 | 61.5 | 19.6 |
R50 | 110 | 68.3 | 68.1 | 69.7 | 68.7 | 19.6 | |
StreamMapNet [31] | R50 | 24 | 64.1 | 58.2 | 59.4 | 60.6 | 19.9 |
MapQR(Ours) | R50 | 24 | 68 | 63.4 | 67.7 | 66.4 | 17.9 |
R50 | 110 | 74.4 | 70.1 | 73.2 | 72.6 | 17.9 | |
swinT | 24 | 68.1 | 63.1 | 67.1 | 66.1 | 14.8 | |
swinB | 24 | 72.1 | 67.5 | 70.5 | 70.1 | 8.3 |
Method | Dim | APdiv | APped | APbou | mAP1 | APdiv | APped | APbou | mAP2 |
{0.2m, 0.5m, 1.0m} | {0.5m, 1.0m, 1.5m} | ||||||||
PivotNet [6] | 2 | 47.5 | 31.3 | 43.4 | 40.7 | - | - | - | - |
MapTR [14] | 40.5 | 27.9 | 32.9 | 33.7 | 58.7 | 55.4 | 59.1 | 57.8 | |
MapTRv2 [15] | 53.0 | 34.2 | 38.8 | 42.0 | 72.1 | 62.9 | 67.1 | 67.4 | |
MapQR | 56.3 | 36.5 | 42.5 | 45.1 | 72.3 | 64.3 | 68.1 | 68.2 | |
MapTRv2 [15] | 3 | 48.1 | 30.5 | 36.7 | 38.4 | 69.1 | 59.8 | 65.3 | 64.7 |
MapQR | 49.4 | 30.6 | 38.9 | 39.6 | 71.2 | 60.1 | 66.2 | 65.9 |
Training and Inference Details. We follow most settings as in the MapTR series [14, 15], and the modifications will be emphasized. The BEV feature is set to be to perceive [] for rear to front, and [] for left to right. We use instance queries to detect map element instances, and each instance is modelled by sequential points. (We set and .)
@@ 4.2 Comparisons with State-of-the-art Methods
Results on nuScenes. Following the previous experiment settings, 2D vectorized map elements are predicted by different methods. The experimental results are listed in Tab. 1 and Tab. 2. It can be seen that the proposed MapQR outperforms all other SOTA methods by a large margin for both mAP1 and mAP2, under the same settings (i.e., backbone and training epoch). Compared with MapTR and MapTRv2, BeMapNet [22] and PivotNet [6] get better results under smaller thresholds (i.e., mAP1) thanks to their elaborate modeling. Still, the proposed method without elaborate modeling for map elements can outperform all of them. It implies that our instance query and improved encoder benefit for more accurate prediction. The inference speed of our method is about FPS, which meets the efficiency requirements of many scenarios.
@@
Results on Argoverse 2. In Tab. 3, we provide the experimental results on Argoverse 2. These testings are all trained epochs with ResNet50 as the backbone. Since Argoverse 2 provides 3D vectorized map elements as ground truth, 3D map elements can be predicted directly () as in MapTRv2. For both 2D and 3D prediction, the proposed MapQR achieves the best performance, especially for stricter thresholds (mAP1). More detailed experimental results are included in the supplementary material.
Method | mAP1 | mAP2 | FPS |
MapTR | 27.3 | 50.3 | 24.2 |
MapTR+SGQ | 34.4(↑7.1) | 58.0(↑7.7) | 22.4 |
MapTRv2 | 37.2 | 61.5 | 19.6 |
MapTRv2+SGQ | 40.1(↑2.9) | 63.4(↑1.9) | 19.5 |
StreamMapNet | 36.7 | 61.1 | 19.9 |
StreamMapNet+SGQ | 38.4(↑1.7) | 62.1(↑1.0) | 18.2 |
Scatter & Gather | Pos. Embed. | mAP1 | mAP2 |
- | - | 28.7 | 51.3 |
- | ✓ | 29.9 | 52.6 |
✓ | - | 32.8 | 56.2 |
✓ | ✓ | 34.4 | 58.0 |